A twelve year 5,000 par valued floating rate bond with annual coupons has a price P such that the investor’s APY is 8%. The annual coupon rate for the n-th year is given by the function .07+.0035*n (please no excel formulas)

a. Find the price P

b. Find the amount of Principal P6 in the sixth coupon payment

c. Find the interest I5 in the fifth coupon payment

d. Find the amount of the discount or premium for this bond.
Answer #1
Year Coupon Rate Cash Flow
[5000*coupon rate]
Discounting Factor
[1/(1.08^year)]
Present Value
[cash flows* discounting factor]
Year Discounting Factor
[1/(1.08^year)]
Present Value
[cash flows* discounting factor]
1 0.07+(0.0035* 1)= 0.0735 367.5 0.925925926 340.2777778
2 0.07+(0.0035* 2)= 0.077 385 0.85733882 330.0754458
3 0.07+(0.0035* 3)= 0.0805 402.5 0.793832241 319.517477
4 0.07+(0.0035* 4)= 0.084 420 0.735029853 308.7125382
5 0.07+(0.0035* 5)= 0.0875 437.5 0.680583197 297.7551487
6 0.07+(0.0035* 6)= 0.091 455 0.630169627 286.7271802 1 0.925925926 421.2962963
7 0.07+(0.0035* 7)= 0.0945 472.5 0.583490395 275.6992118 2 0.85733882 405.0925926
8 0.07+(0.0035* 8)= 0.098 490 0.540268885 264.7317534 3 0.793832241 388.9777981
9 0.07+(0.0035* 9)= 0.1015 507.5 0.500248967 253.8763508 4 0.735029853 373.0276503
10 0.07+(0.0035* 10)= 0.105 525 0.463193488 243.1765812 5 0.680583197 357.3061784
11 0.07+(0.0035* 11)= 0.1085 542.5 0.428882859 232.6689512 6 0.630169627 341.8670226
12 0.07+(0.0035* 12)= 0.112 560 0.397113759 222.3837048 7 0.583490395 326.7546213
12 5000 0.397113759 1985.568793 7 0.583490395 2917.451976
Price P =
(Sum of above PVs)
5361.170914 Principal Value in 6th coupon payment
(Sum of above PVs)
5531.774136

I5 = 437.5

Premium = Price P - Par Value = 5361.17-5000 = 361.17

(If this was helpful then please rate positively. Thank You:)